# Expected Points Above Average: A Novel NBA Player Metric Based on Bayesian Hierarchical Modeling

### v1

@@ Abstract
Team and player evaluation in professional sport is extremely important given the financial implications of success/failure. It is especially critical to identify and retain elite shooters in the National Basketball Association (NBA), one of the premier basketball leagues worldwide because the ultimate goal of the game is to score more points than one’s opponent. To this end we propose two novel basketball metrics: “expected points” for team-based comparisons and “expected points above average (EPAA)” as a player-evaluation tool. Both metrics leverage posterior samples from Bayesian hierarchical modeling framework to cluster teams and players based on their shooting propensities and abilities. We illustrate the concepts for the top 100 shot takers over the last decade and offer our metric as an additional metric for evaluating players.
@@ 1 Introduction
Sports analytics research has three primary goals: evaluating individual impact, improving winning, and quantifying components of the game. These pillars often go hand-in-hand; analytical tools quantifying player impact or features of a game can also support changes in game strategy, which may lead to more winning. With these goals in mind, we revisit the classic problem of measuring player contributions for professional basketball players in the National Basketball Association (NBA). Naively, because the objective of basketball is to score more points than the opponent, it may seem that the most valuable players are those who score the most points. However, this approach ignores other aspects of the game, such as the skill of the other players on the team. A better way to evaluate a player might be by their shooting percentage, defined as the number of made shots divided by the total number of shots taken. This approach also falls short since it succumbs to the small sample fallacy, whereby players with few shots have extremely variable make percentages.
@@
There are several sophisticated metrics by which to evaluate players; two popular metrics are Player Efficiency Rating (PER) and Box Plus/Minus (BPM). PER incorporates a variety of basketball statistics for each player’s offensive and defensive contributions (Hollinger, 2007), such as rebounds, assists, and steals. One benefit of PER is it is measured on a per-minute basis, removing any worry about small sample sizes of playing time for a player.
@@
However, a criticism of PER is that it is not interpretable. Alternatively, BPM takes a regression approach and calculates a number defined as the points contributed by a player per 100 possessions above an average player (Myers, 2020). BPM incorporates similar statistics as PER – including both offensive and defensive – and considers a player’s position. While BPM is more interpretable than PER since the value is in points, a big drawback is it relies on a player’s position which is more ambiguously defined in the modern NBA (McMahan, 2018).
@@
Deshpande and Jensen (2016) argue that metrics such as PER and BPM which use statistics like points, rebounds, assists, etc. fail to account for game context. With this motivation, Deshpande and Jensen (2016) propose quantifying the contribution of a player in terms of their team’s odds of winning a basketball game, attempting to control for elements of the game such as “high-leverage" and ‘low-leverage" situations (when the teams are close in score or far away in score, respectively). Alternatively, Daly-Grafstein and Bornn (2020) investigated an individual player’s impact on the game by modeling how a defensive player effects the shot trajectory of an offensive player, ignoring the location of the field goal itself.
With the goals of quantifying aspects of the game and determining the best strategies for winning, other basketball research has implicitly considered the problem of measuring player performance. Sandholtz, Mortensen and Bornn (2020) quantify the efficiency of different groupings of NBA teammates using a Bayesian spatial modeling approach relying on a metric measuring a player’s shot efficiency given which teammates were also on the court. Similarly, Sandholtz and Bornn (2020) used Markov decision processes to analyze shooting strategies by investigating NBA player and shot location data. Jiao, Hu and Yan (2021) used a Bayesian marked spatial point process model to analyze NBA shot charts and clustered players into groups according to their individual shot taking data. Yin, Hu and Shen (2022) assessed the latent heterogeneity in the selection of field goal attempts by NBA players using a mixture of finite mixtures framework, examining clusters of players via Bayesian inference which allows for estimating the optimal number of clusters while also estimating cluster membership.
@@
In this paper we creatively translate a metric for player comparison originally developed for baseball into the basketball arena. We call it expected points above average (EPAA). This metric helps teams better identify players to drive winning. Our derivation of the metric also quantifies player and team shooting profiles in a way which offers insight for comparing teams and players. Our work emulates the runs above average statistic commonly used in Major League Baseball (Winston, 2012) to evaluate players based on their expected point contribution to the average team above what an average player would contribute. A parallel metric for evaluating players is called wins above replacement, where the statistic captures the increase in the number of wins a player contributes to their team as opposed to if they were replaced with a replacement player. A replacement player is usually defined as a player who is somehow accepted as lesser than most (perhaps measured by low levels of playing time). Wins above replacement has been used in baseball, see Baumer, Jensen and Matthews (2015) where the authors used runs above average to calculate wins above replacement. It also has found use in American football, e.g. Yurko, Ventura and Horowitz (2018), and hockey (Thomas and Ventura, 2015). These metrics are nicely interpretable and have widespread use in professional sports. This led to us adopting the same form for our metric which, like Winston (2012), is a measure of the contribution of a player over and above the average player as opposed to that of a replacement player.
@@
We perform both a team- and player-level analysis, whereby we consider the location of shots taken on the basketball court and the shot making propensities for each player or team. Our work is similar to that of Jiao, Hu and Yan (2021) in that we both use Bayesian hierarchical clustering of NBA players (and teams in our case). A key difference, however, is that we utilize discretized shot locations rather than the specific $(x,y)$ coordinates on the court as the basis of our model. We do this because there are different point values associated with different regions on the court, e.g. a (made) left corner three is worth three points whereas a basket in the paint is only worth two. Spatial modeling in the Euclidean plane does not capture these nuances. Furthermore, we are less interested in the resulting cluster assignments of players (and teams) than Jiao, Hu and Yan (2021). We share their belief that players and teams have shooting commonalities, but we use these estimated clusters and their associated shooting characteristics to examine the impact of a specific player relative to an "average" player. We illustrate our data and analysis pipeline, and highlight its novelty, in Figure 1.
@@
Figure 1: A schematic of the analytical pipeline used to the compute the EPAA is shown above. At the top left, we first cluster players using a Bayesian hierarchical model applied to their shot taking and shot making data. We then summarize the player clusters using the posterior distribution of the shot taking probabilities across the six areas of the court and the corresponding make probability for each region (top middle). Simultaneously, we use the Bayesian hierarchical model to cluster teams (bottom left) and estimate team cluster characteristics. From the team cluster shot making/taking probabilities, we derive the posterior distribution of expected points scored in a season. We then average the cluster-specific expected point distributions to obtain the distribution of expected points for an average team (bottom right). Finally, to obtain the EPAA of player X (top right), we use the posterior distribution of the player cluster memberships, and the corresponding cluster characteristics, to compute the expected points of the average team if we substitute in player X. We compare this distribution (shown in red in the top right figure) to the expected point distribution for the average team and define the EPAA as the difference between the means of these distributions.
@@
By using the discretized shot locations, we induce a Multinomial distribution of shot selection across the court parameterized by the total number of shots taken and the propensity to take shots in the various regions. We then apply a Bayesian hierarchical modeling framework to our Multinomial data to identify team or player clusters based on both the shot taking and shot making patterns.
We use the results of this team-based clustering model to derive a team-specific metric called expected points (EP) in order to compare teams within and across seasons. Next, we define a player-specific metric, expected points above average (EPAA), by combining the results of both team and player clustering using the proposed Bayesian Hierarchical framework. EPAA offers a new way to evaluate players by measuring their impact on an “average” NBA team.
@@
This paper is organized as follows. In the next section we describe the data used for the analyses. In Section 3 we describe the statistical model and method of inference. The results of our analysis are summarized in Section 4. Section 5 concludes with a discussion surrounding our main results and directions of future work.
@@ 3.1 Bayesian Hierarchical Model Specification
Within the Bayesian framework we assign probability distributions to the parameters $p_{w}^{k}$ and $q_{z}^{k}$ for all clusters $w$ and $z$. For $w\in\{1,\dots,L\}$, each length-$K$ vector of shot selection parameters is modeled independently as
@@
We use a Gibbs sampling algorithm to sample from the joint posterior distribution of the parameters given team shot counts $\boldsymbol{N}=\big{\{}\{N_{i}^{k}\}_{k=1}^{K}\big{\}}_{i=1}^{I}$ and team shot makes $\boldsymbol{M}=\big{\{}\{M_{i}^{k}\}_{k=1}^{K}\big{\}}_{i=1}^{I}$. The basic mechanics involve sampling sequentially from a series of full conditional distributions to obtain a posterior draw from each parameter in the set $\boldsymbol{\omega}=\{\boldsymbol{w},\boldsymbol{z},\boldsymbol{p},\boldsymbol{q},\boldsymbol{\pi},\boldsymbol{\theta}\}$, where $\boldsymbol{p}$ is a length-$(K\times L)$ vector of shot selection parameters (see (1)), $\boldsymbol{q}$ is a length-$(K\times J)$ vector of shot accuracy parameters (see (2)), $\boldsymbol{w}$ is a length-$I$ vector of shot selection cluster memberships, $\boldsymbol{z}$ is a length-$I$ vector of shot accuracy cluster memberships, and $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{L})^{\prime}$ and $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{J})^{\prime}$ are the vectors of cluster membership probabilities. See Franzén (2006) for a general description of Gibbs sampling in the Bayesian mixture model context.
@@ 3.2.2 Expected Points Above Average
Let $\tilde{N}^{i}$ denote a hypothetical number of shots taken by player $i$. To compute EPAA, we compute the difference in expected points based on $\tilde{N}^{i}$ shots had they been taken by player $i$ versus if they were taken by an “average” team. We refer to “average” in this case to refer to a randomly selected team and the expected points based on that team with $\tilde{N}^{i}$ shots. In both cases (player and average team), we sample from the posterior predictive distributions as described in Section 3.2.1 given $\tilde{N}^{i}$ total shots. For player $i$, we sample from the posterior predictive distribution given $\tilde{N}^{i}$ shots based on the player-level model. The distribution of the difference in point totals between the player and the average team then represents the impact of player $i$.
4 Data Analysis
We initialize the Gibbs sampler in the following manner. We set $L$ and $J$, the number of shot selection and shot accuracy clusters, respectively, to $L=J=10$, which allows for a total of 100 shot selection and shot accuracy cluster combinations. We investigated other choices of $L$ and $J$ (e.g., $L=J=5$; $L=5$, $J=10$; $L=10$, $J=5$) and the results of our analysis are robust to this initial specification. We set $\alpha=\beta=\gamma=5$.
@@ 4 Data Analysis
It should be noted that as part of this analysis we conducted a sensitivity study with respect to the choice of $L$ and $J$, and the hyperparameters, $\alpha$, $\beta$, and $\gamma$. The full results of that sensitivity analysis can be found in Section S1 in this manuscript’s online supplement. The primary conclusion shows that the overall conclusions presented here are only marginally affected by our choices.
@@ 4.2 Expected Points and Expected Points Above Average
We compute the posterior predictive distribution of expected points for each team in the 2020-21 season based on the approach outlined in Section 3.2. The predictive distributions of expected points are displayed in Figure 4 for each team. Based on our model and historical performance we would expect the top scoring team to be the 2020-21 Brooklyn Nets with an estimated total of 120 points per game based on 8000 shots over the season. The 2020-21 Nets were led by Kevin Durant, James Harden, and Kyrie Irving, three players known for their scoring prowess. The Utah Jazz, LA Clippers, and Denver Nuggets round out the top four teams. The bottom four teams are the Minnesota Timberwolves, Oklahoma City Thunder, Cleveland Cavaliers, and the Orlando Magic.
@@
Figure 5 shows posterior mean and 95% credible intervals estimates of EPAA for nineteen players from the 2020-21 season. The list of players includes fourteen of the All NBA first, second, and third teams (Rudy Gobert is omitted due to a lack of shots) as well as five additional players with high EPAA who were not selected for an All NBA team (highlighted in maroon). Note, All NBA teams are a recognition of the top 15 (5 per team) players as voted by members of the sports media community. The five players in maroon could be considered All NBA “snubs”, or players whose success could conceivably have warranted them a spot on All NBA but were omitted.This analysis shows that traditionally high-volume and accurate three point shooters, e.g. Bradley Beal, Stephen Curry, Kyrie Irving, and Damian Lillard, tend to have the largest impact above an average team taking the same number of shots. Note also that these four players play one of the two guard position, either a point or shooting guard. Interestingly, however, the 2020-21 NBA Most Valuable Player, Nikola Jokić, who plays center, is ranked second in terms of mean EPAA. This indicates EPAA can be used as an additional objective measure for a player’s All NBA portfolio.
@@
It may be tempting to attribute the ordering in Figure 5 to the proportion of shots taken by each player relative to their team during that season. However there is no meaningful correlation between a player’s EPAA and the proportion of the team’s shots taken by the player (see Figure 6). In fact, a diagnostic such as this could be used to advocate on behalf of players that are not shooting at a very high volume relative to their team and yet still have a high EPAA (upper left), e.g. Joe Harris and Kendrick Nunn.
@@
Next, we investigate the ranking within year of the posterior predictive mean EPAA of ten of the best players from the 2010s across each season for which they qualified (i.e., one of the top 100 shot takers per season). The ten players were chosen based on a query to ChatGPT asking for the top fifteen players of the 2010s. These players are the top eleven minus Kobe Bryant, who only qualified for three years in the 2010s. The results are given in Figure 7. It is interesting to note that Dirk Nowitzki ranked in the top 12, and very often top five, even though he was nearing the end of his career in the 2010s. Furthermore, it shows how dominant Stephen Curry and Kevin Durant have been through this time frame. We also clearly see the peak and decline of Russell Westbrook and the gradual decline of Dwyane Wade as he ended his career.
@@
Figure 8: The two panels show the mean EPAA for the top one hundred shot takers in the 2020-21 NBA season plotted against: (a) the player’s Box Plus/Minus value and (b) the player’s Player Efficiency Rating for the same season.
5 Web Application, Data, and Code
To increase the utility of EPAA as a player evaluation tool, we provide a Shiny application (Chang et al., 2023) as a means to compare the performance of players across three seasons (2017-18, 2018-19, and 2020-21). The application has five key components: (1) density plots of EPAA for up to four players per season, (2) a sortable table of all players for which EPAA was calculated, (3) a download button with access to the posterior draws for EPAA calculations, (4) an interactive scatterplot showing a team’s shot taking and accuracy trends over time, and (5) EPAA over time for up to four players. The full application can be found at (Elmore, 2023) and the landing page is given in Figure 9.
Figure 9: The Shiny application showing the default EPAA player-comparison density plots for the 2020-21 season.

### v2

@@ Abstract
In this paper, we propose two novel basketball metrics: “expected points” for team-based comparisons and “expected points above average (EPAA)” as a player-evaluation tool. Established within the Bayesian hierarchical model framework, teams and players are clustered based on their shooting propensities and abilities using posterior predictive distributions. We illustrate the concepts for the top 100 shot takers over the last decade and offer our metric as an additional metric for evaluating players. We compare our metrics to two traditional NBA player evaluation metrics: player efficiency rating and box plus/minus. Finally, we develop a Shiny web application that allows interested readers to make additional team and player comparisons.
@@ 1 Introduction
Sports analytics research has three primary goals: evaluating individual impact, improving winning, and quantifying components of the game. These pillars often go hand-in-hand; analytical tools quantifying player impact or features of a game can also support changes in game strategy, which may lead to more winning. With these goals in mind, we revisit the classic problem of measuring player contributions for professional basketball players in the National Basketball Association (NBA). Naively, because the objective of basketball is to score more points than the opponent, it may seem that the most valuable players are those who score the most points. However, this approach ignores other aspects of the game, for example the skill(s) of the other players on the team. A better way to evaluate a player might be by their shooting percentage, defined as the number of made shots divided by the total number of shots taken. This approach also falls short since it succumbs to the small sample fallacy, whereby players with few shots may have extremely variable make percentages.
@@
There are several sophisticated metrics by which to evaluate players; two popular metrics are Player Efficiency Rating (PER) and Box Plus/Minus (BPM). PER incorporates a variety of basketball statistics for the offensive and defensive contributions of each player (Hollinger, 2007), such as rebounds, assists, and steals. One benefit of PER is that it is measured on a per-minute basis, removing any worry about small sample sizes of playing time for a player. However, a criticism of PER is that it is not interpretable.
@@
Alternatively, BPM is a regression-based approach and is defined as the points contributed by a player per 100 possessions above an average player (Myers, 2020). BPM incorporates statistics similar to PER–including both offensive and defensive–and considers the position of the player. Although BPM is more interpretable than PER since its units is in points, a major drawback is that it relies on the position of the player, which is defined more ambiguously in the modern NBA (McMahan, 2018).
@@
Deshpande and Jensen (2016) argue that metrics such as PER and BPM which use statistics such as points, rebounds, assists, etc. fail to account for game context. With this motivation, Deshpande and Jensen (2016) propose quantifying the contribution of a player in terms of their team’s odds of winning a basketball game, attempting to control for elements of the game such as “high-leverage” and ‘low-leverage” situations (when the teams are close in score or far away in score, respectively). Alternatively, Daly-Grafstein and Bornn (2020) investigated an individual player’s impact on the game by modeling how a defensive player effects the shot trajectory of an offensive player, ignoring the location of the field goal itself.
With the goals of quantifying aspects of the game and determining the best strategies for winning, other basketball research has implicitly considered the problem of measuring player performance. Sandholtz, Mortensen and Bornn (2020) quantify the efficiency of different groupings of NBA teammates using a Bayesian spatial modeling approach relying on a metric measuring a player’s shot efficiency given which teammates were also on the court. In an attempt to validate known relationships between winning and player performance, and to potentially identify new relationships, D’Urso, De Giovanni and Vitale (2022) estimated a multivariate Bayesian network to study how different player variables affect winning.
Previous efforts have also looked at shot location data. Sandholtz and Bornn (2020) used Markov decision processes to analyze shooting strategies by investigating NBA player and shot location data. Reich et al. (2006) used hierarchical spatial models for shot-chart data for individual players, with one stated goal being to develop a strategy for guarding specific players.
Turning to clustering players based on similar shooting characteristics, Jiao, Hu and Yan (2021) used a Bayesian marked spatial point process model to analyze NBA shot charts and clustered players into groups according to their individual shot taking data. Hu, Yang and Xue (2021) utilized a mixture of finite mixtures model to examine the heterogeneity of shot locations of players to find players with similar shooting profiles. Yin, Hu and Shen (2022) extended this work by examining clusters of players via Bayesian inference, which allows for estimating the optimal number of clusters while also estimating cluster membership. Clustering was also the goal of Muniz and Flamand (2022) who employed k-Means clustering of player statistics to identify outlying players and player archetypes. Others have gone beyond simply using player statistics to measure player performance, for example Zhang et al. (2018) used both player height and weight measurements in addition to playing statistics for constructing player clusters.
@@
In this paper, we creatively translate a metric for player comparison originally developed for baseball into the basketball arena. We call it expected points above average (EPAA). This metric helps teams better identify players that drive winning. Our derivation of the metric also quantifies player and team shooting profiles in a way which offers insight for comparing teams and players. Our work emulates the runs above average statistic commonly used in Major League Baseball (Winston, Nestler and Pelechrinis, 2022; Elmore and Urbaczewski, 2025) to evaluate players based on their expected point contribution to the average team above what an average player would contribute. A parallel metric for evaluating players is called wins above replacement, where the statistic captures the increase in the number of wins a player contributes to their team as opposed to if they were replaced with a replacement player. Wins above replacement has been used in baseball, see Baumer, Jensen and Matthews (2015) where the authors used runs above average to calculate wins above replacement. It also has found use in American football, e.g. Yurko, Ventura and Horowitz (2018), and hockey (Thomas and Ventura, 2015). These metrics are nicely interpretable and have widespread use in professional sports. This led us to adopt the same form for our metric which is a measure of the contribution of a player above the average player as opposed to that of a replacement player.
@@
We perform both a team- and player-level analysis, whereby we consider the location of shots taken on the basketball court and the shot making propensities for each player or team. Our work is similar to that of Jiao, Hu and Yan (2021) in that we both use Bayesian hierarchical clustering of NBA players (and teams in our case). A key difference, however, is that we utilize discretized shot locations rather than the specific $(x,y)$ coordinates on the court as the basis of our model. We do this because there are different point values associated with different regions on the court, e.g. a (made) left corner three is worth three points whereas a basket in the paint is only worth two. Spatial modeling in the Euclidean plane does not capture these nuances. Furthermore, we are less interested in the resulting cluster assignments of players (and teams). We agree with Jiao, Hu and Yan (2021) that players and teams have shooting commonalities, but we use these estimated clusters and their associated shooting characteristics to examine the impact of a specific player relative to an ”average” player. We illustrate our data and analysis pipeline, and highlight its novelty, in Figure 1.
@@
Figure 1: A visual representation of the analytical pipeline used to the compute the EPAA. Along the top, we illustrate the clustering players based on their shot taking and shot making data (top left) using Bayesian hierarchical modeling. We then summarize the player clusters using the posterior distribution of the shot taking probabilities and the corresponding make probability for each region (top middle). Simultaneously, we use similar modeling framework to cluster teams (bottom left) and estimate team cluster characteristics. From the team cluster shot making/taking probabilities, we derive the posterior distribution of expected points scored in a season (bottom middle). We then average the cluster-specific expected point distributions to obtain the distribution of expected points for an average team (bottom right). Finally, to obtain the EPAA of player X (top right), we use the posterior distribution of the player cluster memberships, and the corresponding cluster characteristics, to compute the expected points of the average team if we substitute in player X. We compare this distribution (shown in red in the top right figure) to the expected point distribution for the average team and define the EPAA as the difference between the means of these distributions.
@@
By using the discretized shot locations, we induce a multinomial distribution of shot selection across the court parameterized by the total number of shots taken and the propensity to take shots in the various regions. We then apply a Bayesian hierarchical modeling framework to our multinomial data to identify team or player clusters based on both the shot taking and shot making patterns. We use the results of this team-based clustering model to derive a team-specific metric called expected points (EP) to compare teams within and across seasons. Next, we define a player-specific metric, expected points above average (EPAA), by combining the results of both team and player clustering using the propose Bayesian Hierarchical framework. EPAA offers a new way to evaluate players by measuring their impact on an “average” NBA team. See Section 3.2.2 for more details related to how we define “average” in this research.
The Bayesian model employed in our analysis has been studied extensively in the past (see, for example, Nandram, 1998). As such, the novelty of our work lies in the analytical pipeline described in Figure 1. Using posterior sampling, an important deliverable of our approach is the creation of new basketball player and team evaluation tools.
@@
The remainder of the paper is organized as follows. The next section describes the data used for the analyses. In Section 3 we describe the statistical model and method of inference. The results of our analysis are summarized in Section 4. Section 5 concludes with a discussion surrounding our main results and directions of future work.
@@ 3.1 Bayesian Hierarchical Model Specification
[…]
@@
[…]
@@ 3.2.2 Expected Points Above Average
Let $\tilde{N}^{i}$ denote a hypothetical number of shots taken by player $i$. To compute EPAA, we compute the difference in expected points based on $\tilde{N}^{i}$ shots had they been taken by player $i$ versus if they were taken by an “average” team. In both cases (player and average team), we sample from the posterior predictive distributions as described in Section 3.2.1 given $\tilde{N}^{i}$ total shots.
Specifically, for player $i$, we sample from the posterior predictive distribution given $\tilde{N}^{i}$ shots based on the player-level model. We then sample from the posterior predictive distribution given $\tilde{N}^{i}$ shots using the team-level model. That is, we randomly select a team and use that team’s characteristics for each sample from the team-level model to construct an ”average” team’s distribution of points given $\tilde{N}^{i}$ shots. The distribution of the difference in point totals between the player and the average team then represents the impact of player $i$.
It should be noted that we did not compute the ”average” team based on randomly sampling individual players. This approach would have led to biased results as our individual player model was fitted only to the top 100 shot takers during the 2020-2021 season.
4 NBA Player and Team Data Analyses
We fix the model cluster specification in the following manner. We set $L$ and $J$, the number of shot selection and shot accuracy clusters, respectively, to $L=J=20$, which allows for a total of 400 shot selection and shot accuracy cluster combinations for both players and teams. Furthermore, we set the nuisance parameters at $\alpha=\beta=\gamma=5$ in this analysis. We examined the similarity of the resulting clusters with respect to the estimated shot taking and making probabilities in Section S1 of the online supplement (Williams et al., 2025).
Through a sensitivity analysis, we investigate additional choices of $L$ and $J$, including $L,J\in\{10,20,30\}$ as well as other specifications of the parameters $\alpha=\beta=\gamma$. Based on our exploration, there were only minor differences in expected points per game and EPAA, respectively, when increasing $L$ or $J$ beyond 20 (Figures 3 and 4 and Figures 7 and 8 of the online supplement). We did not detect much change in the expected points per game or EPAA when varying $\alpha$, $\beta$, and $\gamma$ (Figures 5, 6, 9, and 10 in the online supplement). Detailed results of this sensitivity analysis are provided in Section S2 of the online supplement.
@@ 4.2 Expected Points and Expected Points Above Average
We compute the posterior predictive distribution of expected points for each team in the 2020-21 season based on the approach outlined in Section 3.2. The predictive distributions of expected points are displayed in Figure 4 for each team. Based on our model and historical performance we would expect the top scoring team to be the 2020-21 Brooklyn Nets with an estimated total of 120 points per game based on 8000 shots over the season. The 2020-21 Nets were led by Kevin Durant, James Harden, and Kyrie Irving, three players known for their scoring prowess. The LA Clippers, Phoenix Suns, and Denver Nuggets round out the top four teams. The bottom four teams are the New York Knicks, Oklahoma City Thunder, Cleveland Cavaliers, and the Orlando Magic.
@@
Figure 5 shows posterior mean and 95% credible intervals estimates of EPAA for nineteen players from the 2020-21 season. The list of players includes fourteen of the All NBA first, second, and third teams (Rudy Gobert is omitted due to a lack of shots) as well as five additional players with high EPAA who were not selected for an All NBA team (highlighted in maroon). Note, the All-NBA teams are a recognition of the top 15 (5 per team) players as voted by members of the sports media community. The five players in maroon could be considered All NBA “snubs”, or players whose shooting success could conceivably have warranted them a spot on All NBA, but were omitted. This analysis shows that traditionally high-volume and accurate three-point shooters tend to have the largest impact above an average team taking the same number of shots with Stephen Curry ranking first. Note also that Tobias Harris, Michael Porter Jr., and Khris Middleton rank four, five, and six among these players and were not selected for All NBA. Interestingly, however, the 2020-21 NBA Most Valuable Player, Nikola Jokić, who plays center, is ranked second in terms of mean EPAA. This analysis indicates EPAA can be used as an additional objective measure for a player’s All NBA portfolio.
@@
It may be tempting to attribute the ordering in Figure 5 to the proportion of shots taken by each player relative to their team during that season. However, there is no meaningful correlation between a player’s EPAA and the proportion of the team’s shots taken by the player (see Figure 6). In fact, a diagnostic such as this could be used to advocate on behalf of players that are not shooting at a very high volume relative to their team and yet still have a high EPAA (upper left), e.g. Joe Harris, Kendrick Nunn, Mikal Bridges, and Duncan Robinson.
@@
Next, we investigate the ranking within year of the posterior predictive mean EPAA of ten of the best players from the 2010s across each season for which they qualified (i.e., one of the top 100 shot takers per season). The ten players were chosen based on a query to ChatGPT asking for the top fifteen players of the 2010s. These players are the top eleven minus Kobe Bryant, who only qualified for three years in the 2010s. The results are given in Figure 7. It is interesting to note that Dirk Nowitzki ranked in the top 20 (except 2013) even though he was nearing the end of his career in the 2010s. Furthermore, it shows how dominant Stephen Curry and Kevin Durant have been during this time period. We also clearly see the peak and decline of Russell Westbrook and the gradual decline of Dwyane Wade as he ended his career.
@@
Figure 8: The two panels show the mean EPAA for the top one hundred shot takers in the 2020-21 NBA season plotted against: (left) the player’s Box Plus/Minus value and (right) the player’s Player Efficiency Rating for the same season.
5 Web Application
To increase the utility of EPAA as a player evaluation tool, we provide a Shiny application (Chang et al., 2023) as a means to compare the performance of players across three seasons (2017-18, 2018-19, and 2020-21). The application has five key components: (1) density plots of EPAA for up to four players per season, (2) a sortable table of all players for which EPAA was calculated, (3) a download button with access to the posterior draws for EPAA calculations, (4) an interactive scatterplot showing a team’s shot taking and accuracy trends over time, and (5) EPAA over time for up to four players. The full application can be found at (Elmore, 2024) and the landing page is shown in Figure 9. See the online supplement S5 for additional information related to the code to produce the application and the application itself.
Figure 9: The Shiny application showing the EPAA player-comparison density plots from the 2020-21 season for LeBron James, Giannis Antetokounmpo, Stephen Curry, and Russel Westbrook.